LASERSPECTROSCOPY OF THE B^{2}Σ STATE OF LiAr

\underline{R. BRÜHL}, Department of Physics, Penn State University, University Park, PA 16804; D. ZIMMERMANN, Fachbereich Physik, Technische Universität Berlin, Germany.

The rovibrational structure of the transition B^{2}Σ ← X^{2}Σ of LiAr has been investigated by means of high-resolution laser spectroscopy using a supersonic beam for production of the molecules ^{a}. About 100 absorption lines could successfully be assigned for each isotopomer ^{7}LiAr and ^{6}LiAr providing in addition the vibrational numbering of the observed levels v = 0 …3 of the B^{2}Σ state. The analysis of the experimental observations was complicated by the occurrence of perturbations between rovibrational levels of the B^{2}Σ and of the A^{2}Π state which have not been taken into account, up to now. For the B^{2}Σ interaction potential we used a Morse–van der Waals potential as an analytical expression reproducing the observed energy levels with an error margin of 0.015cm^{−1}. The van der Waals parameter C_{6} was fixed to its theoretical value. Our preliminary results for the B^{2}Σ state of ^{7}LiAr are R_{e} = 6.24(5) Å and D_{e} = 33.8(2.0)cm^{−1}.

aR. Brühl and D. Zimmermann, Chem.Phys.Lett. 233 (1995) p.455